OpenStudy (anonymous):

To approach a runway, a plane must begin a 7° descent starting from a height of 3 miles above the ground. To the nearest mile, how many miles from the runway is the airplane at the start of this approach?

OpenStudy (anonymous):

tan(7) = Height/Horizontal Distance tan(7) = 5/x x*tan(7) = 5 x = 5/tan(7) x = 5/0.12278456 x = 40.721732 x = 41 (round to the nearest mile) So the plane needs to be about 41 miles from the runway